OpenStudy (anonymous):

A man stands on the roof of a 10.0 m -tall building and throws a rock with a velocity of magnitude 24.0 m/s at an angle of 34.0 degrees above the horizontal. You can ignore air resistance. Calculate the magnitude of the velocity of the rock just before it strikes the ground. & Calculate the horizontal distance from the base of the building to the point where the rock strikes the ground

OpenStudy (anonymous):

 \[\alpha=34\] V=24m/s h=10m g=9.81 time from the building to hmax: \[t=( V \times \sin \alpha)/g\] now hmax= \[V \times \sin \alpha \times t - (g \times t ^{2})/2=hmax\] \[(V^{2} \times \sin^{2}\alpha)/(2\times g)=hmax\] so the horizontal distance: \[x=V \times \cos \alpha \times(V \times \sin \alpha )/g + V \times \sqrt{(2 \times (h+hmax)/g)}\] \[x \approx48,59 m\] velocity before strike: \[v =\sqrt{g \times \sqrt{(2 \times (h+hmax)/g)} + V^{2} \times \cos^{2} \alpha}\] \[v \approx20,37m/s\] I could make a mistake in calculations sorry for quality i'm newbe